OpenStudy (anonymous):

help! Find an equation equivalent to r =5/1+cos theta in rectangular coordinates.

OpenStudy (jhannybean):

is it \[r=\frac{5}{1+\cos(\theta)}\]

OpenStudy (anonymous):

ok with want

OpenStudy (anonymous):

If you are starting out with, you'd first do: \[ r=\frac{5}{1+\cos\theta} \implies r+r\cos\theta=5\implies \sqrt{y^2+x^2}+x=5 \]